Find the final amount in the following retirement account, in which the rate of return on the account and the regular contribution change over time. $308 per month invested at 6%, compounded monthly, for 3 years; then $492 per month invested at 7%, compounded monthly, for 3 years.
Answer #1

Part 1 :

Finding the future value of the payments after the first three years

Using financial calculator

Input:

PMT = -308

I/Y = 6/12 = 0.5

N = 3*12 =36

Solve for Fv as 12,115.52

PART 2: finding the future value of the payments after the next three years

Using financial calculator

Input:

PV = - 12,115.52

PMT = -492

N = 36

I/Y = 7/12

Solve for FV as 34,583.14

Final amount in the retirement account= $34,583.14